OpenStudy (anonymous):

an unbiases coin is tossed five times. Probability it lands on heads exactly once?

OpenStudy (anonymous):

\[\binom{5}{1}\left(\frac{1}{2}\right)^5\] or just \[5\times \left(\frac{1}{2}\right)^5\]